# Effect of Overheat and Direct Flow Loading on the MEMS Bistable Flow Sensor

### v1

@@ 2.5.1 Beam equilibrium
To understand the sensor’s behavior and interpret our experimental findings, we explore a reduced-order model of the double-clamped heated microbeam with initial curvature, as shown in Figure 1. While previous models in medina2014experimental ; krylov2018IEEE addressed specific aspects of sensor phenomena, such as aerodynamic loading, overheat ratio, air cooling, or dynamic effects, each model focused on distinct aspects. In this work, we present a comprehensive model that describes the combined influence of all these effects and their interactions.
The motion of the beam subjected to an axial force $N$ (positive in compression) and a distributed actuating force $f_{a}$ can be described by the following equation
@@
Here, $z(x,t)$ and $z_{0}(x)$ are the deformed and the initial (as-designed) elevations of the beam above its anchored ends, respectively (see Figure 1(b)); $EI_{yy}$ and $EA$ are the bending and axial stiffness of the beam, where $I_{yy}={bd^{3}}/{12}$ is the second moment of area of the cross-section, $A=bd$ is the cross-sectional area of the beam, $c$ is the viscous damping coefficient, $E$ is the Young’s modulus, $\rho$ is a mass density and $\dot{(\;)}$, $(\;)^{\prime}$ denote the derivative with respect to the time and the axial coordinate $x$, respectively.
The axial force
@@
is generated by both the beam heating of the beam gerson2010electrothermal and by a residual stress $\sigma_{res}$, estimated to be a few MPa krylov2018IEEE ; kessler2018JMEMS . In Eq. (2), $\alpha$ is the thermal expansion coefficient of Si, and $\overline{\theta}$ is the length-averaged difference between the beam temperature $T({x})$ and the ambient temperature $T_{\infty}$. In our case, we use Joule heating by applying a voltage $V_{\mathrm{h}}$ between the ends of the beam to induce axial loading.
The beam is actuated by an electrostatic force, which is parameterized by the voltage $V_{\mathrm{ES}}$, and can be described by the following equation
@@
where $\phi({x})=1/2[1-\cos(2\pi x/L)]$ is the base function, $q$ and $h$ are the deformed and the initial midpoint elevations of the beam, respectively. Note that $h$ is positive and the minus sign is due to the direction of the positive $z$-axis defined in Figure 1b. Substituting the representation Eq. (4) into Eq. (1), multiplying the resulting expression by $\phi(x)$, integrating by parts, and using the orthogonality of base functions in Galerkin’s formalism, we obtain the following equation expressed with non-dimensional variables
@@
For the case of quasi-static actuation, we can omit the terms containing time derivatives, then Eq. (5) could be solved in respect of $w_{m}(V_{\mathrm{ES}})$.
Equation (5) helps to highlight the flow sensor operational scenario. When the device is exposed to an airflow, the temperature of the beam changes, changing the average temperature difference between the beam and the air, $\overline{\theta}$. This leads to a variation in the axial force ($n$ in Table 2) and alters the critical ST and SB voltages expressed through $\beta$. To predict the $V_{\mathrm{ST}}$ and $V_{\mathrm{SB}}$ values and the flow, it is necessary to quantify the convective heat transfer and relate $\overline{\theta}$ to flow velocity over the beam.
For simplicity of analytical prediction, we limit our analysis to steady-state heat transfer, and neglect radiation. In this simplified case, the heat transfer equation in the beam is governed by the equation
@@
where $\kappa$ is the thermal conductivity, $\tilde{h}$ is the average convective heat transfer coefficient, and $P=b+2d$ is the perimeter of the beam’s cross-section. $J=I/A=V_{\mathrm{h}}/(\rho_{e}L)$ and $\rho_{e}$ represent the electric current density and the beam’s material resistivity, respectively. Assuming that $\kappa$ and $\rho_{e}$ are temperature-independent around the operating point, we can find the averaged solution of Eq. (6) as follows
@@
[…]
@@
In this study, we revisit the conclusions of the works kessler2018JMEMS ; krylov2018IEEE ; kessler2020flow ; kessler2021sampling and perform detailed experimental investigations of the influence of several additional, not addressed in the previous works, effects on the sensor. In the case of perpendicular (to the beam’s axis) flow, the drag force was found in krylov2018IEEE to influence the beam behavior. In the present work, the flow direction is varied, but most of the experimental results are obtained for the flow parallel to the beam axis ($\alpha=0^{\circ}$), where the contribution of the drag force is negligible. Therefore, we have not included the corresponding term into the analytical model.
@@ 2.5.2 Frequency response
where $\omega_{0}=4\pi^{2}\sqrt{EI_{yy}/(3\rho AL^{4})}$ is the fundamental frequency approximation of the initially straight unforced beam and $w_{m}(\beta,\overline{\theta})$ is the solution of the equilibrium equation Eq. (5). More information about this derivation can be found in krylov2010dynamic ; krakover2016displacement ; kessler2021sampling .
@@ 3 Results and discussion
In this section, we outline the results of measurements revealing various factors on sensor behavior during operation, such as the overheat ratio, flow loading, noise and vibrations. Finally, for the first time, we demonstrate the real-time beam shape during ST-SB transitions.
@@ 3.1 Joule heating effect
We present first the quantitative outcomes from various runs conducted at different Joule heating voltages ($V_{\mathrm{h}}$=0, 1, 1.33, and 1.66 V). In Figure 5(a, b), we illustrate the variations in $V_{\mathrm{ST}}$ and $V_{\mathrm{SB}}$ in response to Joule heating at different $V_{\mathrm{h}}$ values across a flow velocity range of 0 to 15 m/s. Notably, our results show that Joule heating induces a consistent offset in all curves. Interestingly, we find that sensitivity remains consistently high across all voltage levels, reflecting the robustness of the sensor’s performance. Note that for all these measurements, we used LDV with the inherent bias in $V_{\mathrm{ST}}$ and $V_{\mathrm{SB}}$ due to laser heating of the beam.
@@
To provide a more comprehensive understanding of the impact of Joule heating voltage playing a key role in determining the flow sensor’s overheat ratio, we thoroughly investigate the beam’s response to a range of electrostatic voltages ($V_{\mathrm{ES}}$). We employ an analytical model, which is detailed in Sect. 2.5, to systematically compare the model’s predictions and the experimental results. This in-depth analysis allows us to assess how varying electrostatic voltages influence the sensor’s behavior, shedding light on its performance in different operational scenarios.
@@
Figure 6: (a) Modeled beam response to the electrostatic voltage $V_{\mathrm{ES}}$ at various heating voltages $V_{\mathrm{h}}$, with a few curves shown for clarity. Insets show the convex (before ST) and concave (after ST) configurations of the beam. (b) Comparison between experimental (symbols) and modeled (solid curve) voltage $\tilde{V}$ as a function of voltage $V_{\mathrm{h}}$. Error bars represent a standard error of the mean. (c) The critical voltage indicator ($\tilde{V}$) as a function of the flow velocity $U$ at different $V_{\mathrm{h}}$ levels. The model results are represented by lines (dotted, dash-dot, dashed, and solid for $V_{\mathrm{h}}$=0, 1, 1.33, and 1.66 V, respectively). Error bars represent a standard error.
@@
As $V_{\mathrm{h}}$ increases, $V_{\text{ST}}$ increases, while the $V_{\text{SB}}$ decreases. Once a certain level of heating is reached, $V_{\text{SB}}$ reaches zero. The meaning is that the beam can stay in the switched configuration at zero actuating voltage. At even higher $V_{\mathrm{h}}$ , a force in the opposite (negative $z$) direction should be applied to switch the beam back. This behavior is known as the latching effect simitses2012dynamic .
To quantify the relationship between critical voltages $V_{\text{ST}}$ and $V_{\text{SB}}$, with respect to both flow velocity and Joule heating, we introduce a universal critical voltage indicator denoted as $\tilde{V}$, calculated as $\tilde{V}=(V_{\text{ST}}-V_{\text{SB}})/(V_{\text{ST}}+V_{\text{SB}})$.
This introduction of a universal indicator serves two key purposes. First, it consolidates the information into a single value, streamlining the analysis rather than dealing with separate ST and SB voltage values. Second, it attempts to disentangle the influence of the overheat ratio effect from the impact of varying ambient temperatures, a factor that can fluctuate in different experimental conditions.
Figure 6(b) compares the analytical model with the experimental results, illustrating the changes in $\tilde{V}$ as a function of $V_{\mathrm{h}}$. The model predictions are consistent with the experimental observations. Furthermore, it predicts the value of the heating voltage value ($V_{\text{h}}=2.2$ V) corresponding to latching, which was not reached in the experiment.
The Joule heating effect in terms of $\tilde{V}$ as a function of the flow velocity $U$ is shown in Figure 6(c). The figure also shows, represented by the different types of lines, the theoretical results obtained using Eq. (5). There is no perfect agreement between the model and the experiment, especially for the low heating case ($V_{\mathrm{h}}$=0 and 1 V). Note that the reduced-order model could not take into account the effect of velocity when the Joule heating is “off” ($V_{\mathrm{h}}$=0 V), but it is possible that there is a small temperature difference between the airflow and the microbeam. Nevertheless, the reduced-order model accurately captures the trends in the critical voltage change, indicating that the primary factor in velocity measurement is the change in average beam temperature, as suggested by the model. The graphs also show that all curves have nearly the same slope angle of 0.015 V/(m/s). This suggests that the beam overheat ratio is not a critical parameter for the sensor, which raises the possibility of significantly reducing energy consumption at the required sensitivity level.
@@ 3.2 Flow angle sensitivity
This section delves into the influence of flow angle on the sensor’s response. The sensor’s angle sensitivity is evaluated in the plane parallel to the sensor, as depicted in Figure 2(b, c), with the sensor’s long side aligned with the flow direction. The results of this impact, illustrated in Figure 7, reveal that the $\tilde{V}$ values remain consistent at low velocities. However, as flow velocities increase, deviations appear for different flow directions defined by the angles $\alpha=0^{\circ},15^{\circ}$, and $45^{\circ}$.
Figure 7: Influence of flow direction by changing the angle $\alpha^{\circ}$ ($V_{\mathrm{h}}$=1.33 V). Dashed lines represent visual trends. Error bars represent a standard error of the mean.
The effect of flow direction on the $\tilde{V}$ value is noticeable. Our sensor behaves similarly to miniature hot-wire sensors used for measuring turbulent flows, as shown in reference borisenkov2015multiarray . Due to its simpler manufacturing process and reliable angular response, the bistable microbeam has potential as a component in future multi-array flow sensors, capable of measuring all three flow velocity components.
@@ 3.3 Frequency response and sensitivity to vibrations
In Figure 4, it is evident that the sensors respond to flow velocity even when no LDV or Joule heating is involved. This suggests the existence of an additional mechanism associated with airflow in our experiment. Since the flow runs parallel to the flow sensor, it is unlikely that this effect is attributed to mean flow drag or lift forces, as these forces are negligible in this particular configuration.
We considered that the device might be sensitive to turbulent velocity fluctuations or chip vibrations. To investigate this, we initially determined the natural frequency of the beam’s vibration around various equilibrium configurations using LDV, following a similar approach to our previous study kessler2021sampling .
@@
During the frequency up-sweep, the frequency $f_{a}$ increases linearly with time, ranging from 30 to 120 kHz. We determined the natural frequency of the beam’s vibration, denoted as $f_{n}$, for different equilibrium deflections determined by the DC bias voltage. The measured resonant frequency of the beam, when there is zero electrostatic voltage applied, is approximately 111 kHz.
@@
Figure 8: Frequency of free vibrations around equilibrium as a function of the increasing/decreasing DC bias voltage. The baseline value $f_{0}$ represents the structure’s frequency in the initial unloaded configuration. The bias voltage is normalized by the snap-through voltage $V_{\mathrm{ST}}$. Experimental results for increasing and decreasing voltages are represented by circle and square markers, respectively. The solid and dash-dot lines represent the calculated increasing/decreasing curves.
@@
As expected, the vibration frequency significantly decreases near the ST and SB points. However, it is crucial to note that the experimental data does not reach absolute zero values, which would entail zero frequency and stiffness. This contrasts with the model’s predictions krylov2010dynamic .
The measured frequencies at the ST and SB branches reveal an exceptionally low beam stiffness in close proximity to the critical $V_{\mathrm{ST}}$ and $V_{\mathrm{SB}}$ points. Consequently, even a minor vibration and noise level can induce an ST transition at voltages lower than initially predicted.
We can derive consistent conclusions by examining the beam’s frequency response in the context of electrostatic actuation with a linear voltage ramp. In this analysis, we utilized the same voltage ramp signal, denoted as $V_{\mathrm{ES}}$, as previously detailed in Section 2.3, with a 60-second duration, allowing for a thorough examination of the beam’s vibrations. The LDV signal was sampled at a rate of 1 MHz.
@@
The spectrogram reveals that the fundamental frequency closely mirrors the curves in Figure 8. Notably, a sharp change in the fundamental frequency is evident during the transition between equilibrium states, specifically during snap-through and snap-back. Moreover, we observe the emergence of a higher (second) harmonic vibration mode, surpassing the intensity of the first mode by a factor of 2.75. These results affirm our findings regarding the sensor’s high sensitivity in proximity to the beam’s equilibrium state. They underscore the fact that even a minor level of vibration or noise can trigger snap-through or snap-back transitions. These factors are known to hinder the experimental voltage-frequency dependency curves from reaching the theoretical zero level, as documented in references tilmans1994electrostatically ; hajjaj2017scalable ; torteman2019micro .
Figure 9: Frequency response of the beam vibrations in the form of a spectrogram during the voltage actuation ramp $V_{\mathrm{ES}}$ (60 s). The level of the PSD is normalized to the maximum value. The overheat ratio is $V_{\mathrm{h}}=$0 V. The line that corresponds to $1.8f_{0}$ is a regular noise signal that comes from the voltage amplifier.
@@
As the amplitude of noise increases, its influence becomes apparent in the absolute value of $\tilde{V}$. This provides direct evidence that the beam exhibits remarkable sensitivity to fluctuations, including those induced by the flow, near the critical equilibrium point. Consequently, it is well-suited for sensing turbulent flows. This capability has the potential for applications beyond the measurement of average flow velocity. It can also be used to capture turbulent fluctuations, such as in shear stress measurements, particularly when the sensors are flush with walls. Notably, similar sensors based on hot-wire technology have been developed chew1998dynamic ; sturzebecher2001surface ; ghouila2019unsteady . Our findings open up new possibilities for applying bifurcation-based sensors in this field.

### v2

@@ 2.5.1 Beam equilibrium
To understand the sensor behavior and interpret our experimental findings, we explored a reduced order model of the double-clamped heated microbeam with initial curvature, as shown in Figure 1. Although previous models in [26, 16] addressed specific aspects of sensor phenomena, such as aerodynamic loading, overheat ratio, air cooling, or dynamic effects, each model focused on different aspects. In this work, we present a comprehensive model that describes the combined influence of all these effects and their interactions.
The motion of the beam subjected to an axial force $N$ (positive in compression) and a distributed actuating force $f_{a}$ can be described by the following equation Eq. (1).
@@
Here, $z(x,t)$ and $z_{0}(x)$ are the deformed and the initial (as-designed) elevations of the beam above its anchored ends, respectively (see Figure 1(b)); $EI_{yy}$ and $EA$ are the bending and axial stiffness of the beam, where $I_{yy}={bd^{3}}/{12}$ is the second moment of area of the cross-section, $A=bd$ is the cross-sectional area of the beam, $c$ is the viscous damping coefficient, $E$ is Young’s modulus, $\rho$ is a mass density and $\dot{(\;)}$, $(\;)^{\prime}$ denote the time and the axial coordinate $x$ derivatives, respectively.
The axial force:
@@
is generated by both the beam heating of the beam [27] and by a residual stress $\sigma_{res}$, estimated to be a few MPa [16, 17]. In Eq. (2), $\alpha$ is the thermal expansion coefficient of Si, and $\overline{\theta}$ is the length-averaged difference between the beam temperature $T({x})$ and the ambient temperature $T_{\infty}$. In our case, we use Joule heating by applying a voltage $V_{\mathrm{h}}$ between the ends of the beam to induce axial loading.
The beam is actuated by an electrostatic force, which is parameterized by voltage $V_{\mathrm{ES}}$, and can be described by the following equation
@@
where $\phi({x})=1/2[1-\cos(2\pi x/L)]$ is the base function, $q$ and $h$ are the deformed and the initial midpoint elevations of the beam, respectively. Note that $h$ is positive, and the minus sign is due to the direction of the positive $z$-axis defined in Figure 1(b). Substituting the representation in Eq. (4) into Eq. (1), multiplying the resulting expression by $\phi(x)$, integrating by parts, and using the orthogonality of base functions in Galerkin’s formalism, we obtain the following equation expressed with non-dimensional variables.
@@
For the case of quasi-static actuation, we can omit the terms containing time derivatives, then Eq. (5) could be solved for $w_{m}(V_{\mathrm{ES}})$.
Equation (5) highlights the flow sensor operational scenario: flow reduces the average temperature difference between the beam and the air, $\overline{\theta}$, affecting the axial force ($n$ in Table 2) and altering the critical ST and SB voltages expressed through $\beta$. To predict the $V_{\mathrm{ST}}$ and $V_{\mathrm{SB}}$ values and the flow, it is necessary to quantify the convective heat transfer and relate $\overline{\theta}$ to flow velocity over the beam.
For simplicity of analytical prediction, we limit our analysis to steady-state heat transfer and neglect radiation. In this simplified case, the heat transfer equation in the beam is governed by the equation (6):
@@
where $\kappa$ is the thermal conductivity, $\tilde{h}$ is the average convective heat transfer coefficient, and $P=b+2d$ is the perimeter of the beam’s cross-section. $J=I/A=V_{\mathrm{h}}/(\rho_{e}L)$ and $\rho_{e}$ represent the electric current density and the resistivity of the beam’s material, respectively. Assuming that $\kappa$ and $\rho_{e}$ are temperature-independent around the operating point, we can find the averaged solution of Eq. (6) as follows,
@@
In Eq. (5), $\overline{\theta}$ is given by Eq. (7), and $\tilde{h}$ can be estimated using the convective heat transfer theory. Due to the small size of the beam, the relevant dimensionless parameter $Re=UL/\nu$ is small, and we can use the laminar boundary layer analysis. $\tilde{h}$ can be determined using the non-dimensional Nusselt number, $Nu$, as $\tilde{h}=Nu\,\kappa_{\text{air}}/L$. In our analytical model, we assume that heat exchange is characterized by a laminar boundary layer on a flat plate (microbeam) and utilize the laminar convection coefficient $Nu=0.664\,Pr^{1/3}\,Re^{1/2}$. Here, the nominal values for airflow can be used: Prandtl number, $Pr=0.7$, the thermal conductivity of air ${\kappa_{\text{air}}=0.0257}$ W/(mK), and the kinematic viscosity $\nu=1.5\times 10^{-5}$ m2/s  [16]. The material properties of single crystal silicon and the rest of the constants required for the application of the model are given in Table 3.
@@
In this study, we revisit the conclusions of the works [17, 16, 21, 22] and perform detailed experimental investigations of the influence of several additional, not addressed in the previous works, effects on the sensor.
@@ 2.5.2 Frequency response
where $\omega_{0}=4\pi^{2}\sqrt{EI_{yy}/(3\rho AL^{4})}$ is the fundamental frequency approximation of the initially straight unforced beam and $w_{m}(\beta,\overline{\theta})$ is the solution of the equilibrium equation Eq. (5). More information on this derivation can be found in [29, 30, 22].
@@ 3 Results and discussion
In this section, we outline the results of measurements revealing various factors on sensor behavior during operation, such as the overheat ratio, flow loading, noise, and vibrations. Finally, for the first time, we demonstrate the real-time beam shape during ST-SB transitions.
@@ 3.1 Joule heating effect
We first present the quantitative results of various runs conducted at different Joule heating voltages ($V_{\mathrm{h}}$=0, 1, 1.33, and 1.66 V). In Figure 5(a, b), we illustrate the variations in $V_{\mathrm{ST}}$ and $V_{\mathrm{SB}}$ in response to Joule heating at different $V_{\mathrm{h}}$ values across a flow velocity range of 0 to approximately 15 m/s. Our results demonstrate that Joule heating causes a consistent offset in all curves, and we achieved high sensitivity across all voltage levels, highlighting the sensor’s performance robustness. Note that for all these measurements, we used LDV with the inherent bias in $V_{\mathrm{ST}}$ and $V_{\mathrm{SB}}$ due to laser heating of the beam.
@@
To better understand the flow sensor’s overheat ratio, we studied the beam’s response to various electrostatic voltages that affect Joule heating. ($V_{\mathrm{ES}}$). We employ an analytical model, detailed in Sect. 2.5, to systematically compare the model’s predictions and the experimental results. This in-depth analysis allows us to assess how varying electrostatic voltages influence the sensor’s behavior, shedding light on its performance in different operational scenarios.
@@
Figure 6: (a) Modeled beam response to the electrostatic voltage $V_{\mathrm{ES}}$ at various heating voltages $V_{\mathrm{h}}$, with a few curves shown for clarity. Insets show the beam’s convex (before ST) and concave (after ST) configurations. (b) Comparison between experimental (symbols) and modeled (solid curve) voltage $\tilde{V}$ as a function of voltage $V_{\mathrm{h}}$. Error bars represent a standard error of the mean. (c) Critical voltage indicator ($\tilde{V}$) as a function of flow velocity $U$ at different $V_{\mathrm{h}}$ levels. The model results are represented by lines (dotted, dash-dot, dashed, and solid for $V_{\mathrm{h}}$=0, 1, 1.33, and 1.66 V, respectively). The error bars represent the standard error.
@@
As $V_{\mathrm{h}}$ increases, $V_{\text{ST}}$ increases, while the $V_{\text{SB}}$ decreases. Once a certain level of heating is reached, $V_{\text{SB}}$ reaches zero. The meaning is that the beam can remain in the switched configuration at zero actuating voltage. At even higher $V_{\mathrm{h}}$ , a force in the opposite (negative $z$) direction should be applied to switch the beam back. This behavior is known as the latching effect [31].
To quantify the relationship between critical voltages $V_{\text{ST}}$ and $V_{\text{SB}}$, to both flow velocity and Joule heating, we introduce a universal critical voltage indicator denoted as $\tilde{V}$, calculated as ${\tilde{V}=(V_{\text{ST}}-V_{\text{SB}})/(V_{\text{ST}}+V_{\text{SB}})}$. This introduction of a universal indicator serves two key purposes. First, it consolidates the information into a single value, streamlining the analysis rather than dealing with separate values for ST and SB voltages. Secondly, it attempts to disentangle the influence of the overheating ratio effect from the impact of varying ambient temperatures, a factor that can fluctuate under different experimental conditions. The point is that the sum of the ST and SB voltages, at least within the single-degree-of-freedom model framework, is much less influenced by varying ambient temperatures. Specifically, following the single-degree-of-freedom model the sum $V_{\mathrm{ST}}+V_{\mathrm{SB}}=\mathrm{const}$, and is independent of the temperature (in the case of an actuation by a ”mechanical”, deflection-independent force). Therefore, the indicator is considered a convenient quantity for comparative studies.
Figure 6(b) compares the analytical model with the experimental results, illustrating the changes in $\tilde{V}$ as a function of $V_{\mathrm{h}}$. The model predictions are consistent with the experimental observations. Furthermore, it predicts the value of the heating voltage value ($V_{\text{h}}=2.2$ V) corresponding to latching, which was not reached in the experiment. It is seen that if the latching effect is achieved, the parameter $\tilde{V}$ equals one.
The Joule heating effect in terms of $\tilde{V}$ as a function of the flow velocity $U$ is shown in Figure 6(c). The figure also shows, represented by the different types of lines, the theoretical results obtained using Eq. (5). There is no perfect agreement between the model and the experiment, especially for the low heating case ($V_{\mathrm{h}}$=0 and 1 V). Note that the reduced-order model could not take into account the effect of velocity when the Joule heating is “off” ($V_{\mathrm{h}}$=0 V), but there may be a small temperature difference between the airflow and the microbeam. Nevertheless, the reduced-order model accurately captures the trends in the critical voltage change, indicating that the primary factor in velocity measurement is the change in average beam temperature, as suggested by the model. The graphs also show that all curves have nearly the same slope angle of 0.015 V/(m/s). This suggests that the beam overheat ratio is not a critical parameter for the sensor, which raises the possibility of significantly reducing energy consumption at the required sensitivity level.
@@ 3.2 Flow angle sensitivity
[…]
Figure 7: Influence of flow direction by changing the angle of attack ($V_{\mathrm{h}}$=1.33 V).
Dashed lines represent visual trends. Error bars represent a standard error of the mean.
The effect of flow direction on the $\tilde{V}$ value is noticeable. Our sensor behaves similarly to miniature hot-wire sensors for measuring turbulent flows, as shown in reference [32]. Due to its more straightforward manufacturing process and reliable angular response, the bistable microbeam has potential in future multi-array flow sensors capable of measuring all three flow velocity components.
@@ 3.3 Frequency response and sensitivity to vibrations
In Figure 4, it is evident that the sensors respond to flow velocity even when no LDV or Joule heating is involved. This suggests an additional mechanism associated with airflow in our experiment. Since the flow runs parallel to the flow sensor, this effect is unlikely to be attributed to mean flow drag or lift forces, as these forces are negligible in this particular configuration.
We considered that the device might be sensitive to turbulent velocity fluctuations or chip vibrations. To investigate this, we initially determined the natural frequency of the vibration of the beam around various equilibrium configurations using LDV, following a similar approach to our previous study [22].
@@
During the frequency up-sweep, the frequency $f_{a}$ increases linearly with time, ranging from 30 to 120 kHz. We determined the natural frequency of the beam’s vibration, denoted as $f_{n}$, for different equilibrium deflections determined by the DC bias voltage. The measured resonant frequency of the beam, when zero electrostatic voltage is applied, is approximately 111 kHz.
@@
Figure 8: Frequency of free vibrations around equilibrium as a function of the increasing/decreasing DC bias voltage. The baseline value $f_{0}$ represents the structure’s frequency in the initial unloaded configuration. The snap-through voltage $V_{\mathrm{ST}}$ normalized the bias voltage. Experimental results for increasing and decreasing voltages are represented by circle and square markers, respectively. The solid and dash-dot lines represent the calculated (model) increasing/decreasing curves.
@@
As expected, the vibration frequency decreases near the ST and SB points. However, it is crucial to note that the experimental data does not reach absolute zero values, which would entail zero frequency and stiffness. This contrasts with the model’s predictions [29].
The measured frequencies at the ST and SB branches reveal an exceptionally low beam stiffness near the critical $V_{\mathrm{ST}}$ and $V_{\mathrm{SB}}$ points. Consequently, even a minor vibration and noise level can induce an ST transition at voltages lower than initially predicted.
We can derive consistent conclusions by examining the beam’s frequency response in electrostatic actuation with a linear voltage ramp. In this analysis, we utilized the same voltage ramp signal, denoted as $V_{\mathrm{ES}}$, as previously detailed in Section 2.3, with a 60-second duration, allowing for a thorough examination of the beam’s vibrations. The LDV signal was sampled at a rate of 1 MHz.
@@
The spectrogram reveals that the fundamental frequency closely mirrors the curves in Figure 8. Notably, a sharp change in the fundamental frequency is evident during the transition between equilibrium states, specifically during snap-through and snap-back. Moreover, we observe the emergence of a higher (second) harmonic vibration mode, surpassing the intensity of the first mode by a factor of 2.75. These results affirm our findings regarding the sensor’s high sensitivity in proximity to the beam’s equilibrium state. They underscore that even a minor level of vibration or noise can trigger snap-through or snap-back transitions. These factors are known to hinder the experimental voltage-frequency dependency curves from reaching the theoretical zero level, as documented in references [33, 34, 35].
Figure 9: Frequency response of the beam vibrations in the form of a spectrogram during the voltage actuation ramp $V_{\mathrm{ES}}$ (60 s). The level of the PSD is normalized to the maximum value. The overheat ratio is $V_{\mathrm{h}}=0$ V. The line corresponding to $1.8f_{0}$ is a regular noise signal from the voltage amplifier.
@@
As the amplitude of noise increases, its influence becomes apparent in the absolute value of $\tilde{V}$. This provides direct evidence that the beam exhibits remarkable sensitivity to fluctuations, including those induced by the flow, near the critical equilibrium point. Consequently, it is well-suited for sensing turbulent flows. This capability has the potential for applications beyond the measurement of average flow velocity. It can also capture turbulent fluctuations, such as in shear stress measurements, particularly when the sensors are flush with walls. Notably, similar sensors based on hot-wire technology have been developed [36, 37, 38]. Our findings open up new possibilities for applying bifurcation-based sensors in this field. Note that since LDV sensing is suitable only in a laboratory environment, snap-through events detection based on electrostatic, piezoelectric, or piezoresistive sensing should be developed in the future to be implemented in integrated packaged sensors used in field experiments.